# A causal random synthetic turbulent velocity field: I. The underlying Gaussian structure

### v1

@@ 3.3 A second comparison concerning the temporal structure of DNS and the Gaussian model
We begin by estimating on the DNS field the time correlation of Fourier modes $\widehat{\mathbf{u}}^{\text{\tiny{DNS}},\nu}(t,\mathbf{k}_{\mathbf{n}})$ for several values of the wave vector $\mathbf{k}_{\mathbf{n}}$, as it has been already done by Gorbunova et al. (2021). To be more precise, and because we have downloaded data only in a limited region of space, we will be focusing on the correlation in time of the one-dimensional longitudinal Fourier modes. We derive in Appendix B the precise relationship between the time-correlation structure on the one-dimensional longitudinal Fourier modes
@@
The solid lines of Fig. 2(b) are the theoretical predictions of $F^{(N)}_{\text{\tiny{long}},k_{n}}(\tau)$ for the above mentioned $k_{n}$, computed in Appendix B
@@
Figure 2: Temporal correlation structure of the longitudinal velocity Fourier modes (112). (a) Longitudinal correlation function $C_{\text{\tiny{long}}}^{\nu}(\tau,k_{n})$ (111) of the DNS velocity field $\widehat{u}_{\text{\tiny{long}}}^{\text{\tiny{DNS}},\nu}(t,k_{n})$, for $k_{n}L_{\text{\tiny{tot}}}=42,\ 75,\ 107,\ 141,\ 173,\ 205$. The characteristic large time scale $T_{E}^{\text{\tiny{DNS}}}=1.99$ is defined in the readme file of the DNS. Inset: Re-scaled correlation functions for the same wave numbers $k_{n}$, with $D_{3}$ provided in the text. (b) Similar representation of the longitudinal correlation function $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ (111) as in (a) but for the Gaussian model $\widehat{u}_{\text{\tiny{long}}}^{(2),\nu}(t,k_{n})$. Solid lines represent theoretical predictions while dot-lines are numerical estimations for the same wavenumbers $k_{n}$ as (a). The characteristic large time scale $T_{E}=1.43$ is defined by $T_{E}:=\sqrt{3}\frac{L_{\text{\tiny{int}}}}{\sigma^{\nu}}$ with $L_{\text{\tiny{int}}}$ being the integral time scale defined in (16) Inset: Re-scaled correlation functions for the same wave numbers $k_{n}$, with the same $D_{3}$ as for DNS and provided in the text. (c) Numerical estimation of the characteristic time scale $T_{k_{n}}$ of the time correlation functions displayed in (a) and (b). We superimpose with a dashed black line the theoretical expression (56) with relevant free parameters provided in the text. Inset: re-scaled estimated $T_{k_{n}}$ by $1/(D_{3}k)$, the asymptotical behavior of $T_{k}$ at large $k$.
@@
We show in Fig. 2(a) the results of our numerical estimation of the time correlation function $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ (111) for several wave-number amplitudes $k_{n}$ and as a function of the time-lag $\tau$. At vanishing time lag $\tau=0$, all curves coincide with unity by construction, and then decrease toward 0 (they decorrelate) as $\tau$ increases. The characteristic time scale $T_{k_{\mathbf{n}}}$ of decorrelation is thus clearly a decreasing function of the wave number $k_{n}$. As has been observed by Gorbunova et al. (2021), the very shape of the correlation function is close to a Gaussian function, and so, to estimate the characteristic time scale $T_{k_{n}}$, we perform a similar Gaussian fit as by Gorbunova et al. (2021). In a straightforward manner, for each accessible $1/L_{\text{\tiny{tot}}}<k_{n}<512/L_{\text{\tiny{tot}}}$, we find the best possible $T_{k_{n}}$ such that the temporal correlation function $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ (111) resembles the explicit Gaussian function $\exp(-(\tau/T_{k_{n}})^{2})$ and reproduce the result of this fitting procedure in Fig. 2(c) using an orange line. We can indeed observe that $T_{k_{n}}$ is a decreasing function of the wave number $k_{n}$, and follows a power-law, as has clearly been evidenced by the linear trends in this doubly logarithmic representation. We can also identify two ranges, below and above $k_{n}L_{\text{\tiny{tot}}}\approx 10$. This transition has already been observed in several studies (Kaneda et al., 1999; Favier et al., 2010; Gorbunova et al., 2021) and remains unexplained, as far as we know. We believe that at large scales, i.e. at low wave numbers, the flow is sensitive to boundary conditions and thus cannot be considered universal. At higher wave numbers $k_{n}L_{\text{\tiny{tot}}}$ greater than 10, the power decrease is consistent with the expected behavior given in (56) with $\beta=1/2$ (governing the power-law exponent) and $D_{3}=3.62$. We superimpose with a dashed black line the proposed function dependence of $T_{k_{n}}$ (56) with these aforementioned free parameters and clearly reproduce the decreasing behavior. To check whether such a parameterization of the characteristic time scale reproduces the behaviors of Fourier modes time correlation functions, we add in the inset of Fig. 2(a) a representation of $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ (111) as a function of the re-scaled time variable in the limit of large $k_{n}$ $\tau/T_{k_{n}}\underset{k_{n}\rightarrow+\infty}{\sim}\tau D_{3}k_{n}$, and observe that all correlation functions collapse in a satisfactory manner on a single curve that reads $\exp(-\left(\tau D_{3}k_{n}\right)^{2})$.
@@
For comparison against our modeling approach, we consider the Gaussian velocity field $u^{(2),\nu}_{i}(t,\mathbf{x})$, i.e. for $N=2$. It is fully characterized in the statistically stationary regime by the covariance structure of its Fourier modes $\widehat{u}^{(2),\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})$ which is provided in (95). We have considered an evolution based on two layers $N=2$ that ensures a smooth behavior in time for a finite viscosity, as it is shown in (3.2.4). This choice is dictated by the necessity of ensuring smooth behavior in time of the velocity field, corresponding to bounded velocity time-derivatives (in this case, $\partial_{t}u^{(2),\nu}_{i}(t,\mathbf{x})$ have a finite variance), as is observed in DNS. The case of two embedded layers $N=2$ allows this and is computationally more treatable. As we explain in Chapter 4, each layer needs to be stored in memory, which can become prohibitive given the size of the boxes, which corresponds to three-components of size $1024^{3}$ collocation points for each layer. Besides considering $H=1/3$, the free parameters $D_{2}$, $L$ and $\eta_{d}$ entering in the functional form of the longitudinal PSD (54) have the same values as the ones chosen to describe the spatial structure of the velocity field, as they are given in Section 2.2. The remaining free parameters $D_{3}$ and $\beta$ that appear in the characteristic correlation time scale $T_{k}$ of Fourier modes (56) are the ones that we have estimated before, corresponding to $\beta=1/2$ and $D_{3}=3.62$. We would like to recall that the value of $D_{3}$ is not universal because it is related, when $\beta=1/2$, to the standard deviation of the velocity. An efficient and exact in distribution numerical scheme is proposed in Section 4 in order to give instances in space and time of the model field $u^{(2),\nu}_{i}(t,\mathbf{x})$.
@@
As we did for the DNS velocity field, we display in Fig. 2(b) the temporal longitudinal correlation functions $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ (111) for the modeled Gaussian velocity field $u^{(2),\nu}_{i}(t,\mathbf{x})$. Although being known to be different from the DNS in a theoretical point of view and provided in (95) (once recasted into the longitudinal context (113)), the estimations of $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ for various wave numbers $k_{n}$ displays similar trends as for the DNS. In this case, the collapse is not perfect because we are considering only the longitudinal correlation functions, which necessitate a transformation given in (113)) when going from the three dimensional formulation to the one dimensional setting. We display in Fig. 2(c) the time $T_{k}$ as a function of $k$, for the model for which Fourier mode temporal correlation is prescribed using blue symbols, and the measured Fourier mode correlation time for DNS in orange. We observe a very satisfactory collapse of the estimations of $T_{k}$ from DNS and the model. We also superimpose the asymptotical limit $D_{3}k_{n}$ using the form provided in the inset, we display the re-scaled correlation functions by the asymptotic behavior of the imposed characteristic time scale $D_{3}k_{n}$ (56).
@@
Figure 3: Direct comparison of the temporal structure of the DNS and modeled velocity fields, and estimation of PSDs and second-order structure functions. (a) Spatio-temporal representation of the DNS velocity field $|\mathbf{u}^{\text{\tiny{DNS}},\nu}(t,\mathbf{x})|$ along the spatial line $\mathbf{x}\in([-\pi,\pi],0,0)$ and across time $t\in[0,5027\Delta t]$ where the time stepping $\Delta t$ is provided by the Hopkins database. We use the same characteristic large time scale $T_{E}^{\text{\tiny{DNS}}}$ as in Fig. 2 to adimensionalize time. (b) Similar representation as in panel (a) but for the model $|u^{(2),\nu}_{i}(t,\mathbf{x})|$ with a characteristic large time scale $T_{E}$. We use the same colorbar for panels (a) and (b). (c) Estimation of the temporal power spectral density $E_{\nu}^{\text{\tiny{T}}}(\omega)$ (43) obtained as the variance of the temporal Fourier modes (see text). We use orange for DNS and blue for the model. The dashed black line corresponds to the power-law given in (3.2.4) without any fitting parameters. (d) Estimation of the second-order temporal structure function $S_{2}^{\text{\tiny{T}},\nu}(\tau)$ (given by (2.4) in the inviscid limit), same colors as in (c). Dashed black lines corresponds to two times the variance at large time lags, to (3.2.4) in the inertial range, and to (3.2.4) in the dissipative range.
@@
Our statistical estimation of the temporal power spectrum density $E_{\nu}^{\text{\tiny{T}}}(\omega)$ defined in (43) is shown in Fig. 3(c). To do so, we performed a discrete one-dimensional Fourier transform in time of the profile of the scalar product of velocity fields at a fixed position in space, as required in the definition (43). The obtained evolution in time is not periodic, and in doing so we introduce a spurious discontinuity in the picture. We have checked that windowing this profile aimed at minimizing the consequences of this discontinuity is not altering the observed behaviors in the inertial range of scales, but is slightly changing the results in the respective dissipative range (data not shown). As a result, the temporal spectra correspond to the variance of the (temporal) Fourier modes at each accessible frequency. The variances of the (temporal) Fourier modes of the velocity fields $\mathbf{u}^{\text{\tiny{DNS}},\nu}(t,\mathbf{x})$ and $\mathbf{u}^{(2),\nu}(t,\mathbf{x})$, obtained of the same statistical sample previously described, as functions of the frequency, are displayed doubly logarithmically in Fig. 3(c). We can see that with previously given free parameters, the temporal spectra superimpose in a very satisfactory manner in the inertial range, and follow a typical $\omega^{-5/3}$ power-law decrease. Concerning the model, we observe nevertheless a slight delay before reaching the dissipative range. Let us recall we have two sets of parameters, one entering in the functional form of the longitudinal PSD (54) and the remaining free entering in $T_{k}$ (56). When these parameters are chosen, the temporal PSD shown in Fig. 3(c) is fully determined. We believe that this delay could be weakened while choosing a slightly larger value of the parameter $\eta_{d}$ entering in (54), whose value depends on the simplistic exponential decay used in the present version of the model and/or increasing the number of layers $N$.
@@
We finish with the estimation of the second-order temporal structure function $S_{2}^{\text{\tiny{T}},\nu}(\tau):=\mathbb{E}\left[\left(\mathbf{u}^{\nu}\left(t+\tau,\mathbf{x}\right)-\mathbf{u}^{\nu}\left(t,\mathbf{x}\right)\right)^{2}\right]$, as defined in (2.4) in the vanishing viscosity limit, and display our results in Fig. 3(d). The trends are consistent with the observations made for the temporal spectra. Both DNS and model velocity fields exhibit a power-law behavior in the inertial range, with a typical power-law dependence $\tau^{2/3}$. To see this, we superimpose the prediction made in (3.2.4) with a black dashed line. Note that in the inertial range of scales, predictions depend only weakly on the value of the number of layers N to define the model. Once again, we observe a slight delay of the model to predict the beginning of the dissipative range, for the same reason as the one mentioned in the discussion of the temporal PSDs.
@@
To conclude this comparison, we can say that while the statistical structure of the DNS is fairly satisfactory (at least up to the aforementioned second-order statistical characterization), our model does not capture in a fully appropriate way the advection of the eddies by some large scale flow, as it is clear from the inspection of Fig. 3(a) and Fig. 3(b). To our knowledge, there is no model capable of doing this except some recent propositions by Armstrong & Vicol (2025), where the temporal structure of the advecting velocity field, although simplistic in terms of regularity, is defined recursively through scales using an inverse Lagrangian flow map at each scale of the construction. We will return to these possible perspectives in the final Section.

### v2

@@ 3.3 A second comparison concerning the temporal structure of DNS and the Gaussian model
We begin by estimating on the DNS field the time correlation of Fourier modes $\widehat{\mathbf{u}}^{\text{\tiny{DNS}},\nu}(t,\mathbf{k}_{\mathbf{n}})$ for several values of the wave vector $\mathbf{k}_{\mathbf{n}}$, as it has been already done by Gorbunova et al. (2021). To be more precise, and because we have downloaded data only in a limited region of space, we will be focusing on the correlation in time of the one-dimensional longitudinal Fourier modes. We derive in Appendix C the precise relationship between the time-correlation structure on the one-dimensional longitudinal Fourier modes
@@
The solid lines of Fig. 2(b) are the theoretical predictions of $F^{(N)}_{\text{\tiny{long}},k_{n}}(\tau)$ for the above mentioned $k_{n}$, computed in Appendix C
@@
Figure 2: Temporal correlation structure of the longitudinal velocity Fourier modes (92). (a) Longitudinal correlation function $C_{\text{\tiny{long}}}^{\nu}(\tau,k_{n})$ (91) of the DNS velocity field $\widehat{u}_{\text{\tiny{long}}}^{\text{\tiny{DNS}},\nu}(t,k_{n})$, for $k_{n}L_{\text{\tiny{tot}}}=42,\ 75,\ 107,\ 141,\ 173,\ 205$. The characteristic large time scale $T_{E}^{\text{\tiny{DNS}}}=1.99$ is defined in the readme file of the DNS. Inset: Re-scaled correlation functions in semi-log scale for the same wave numbers $k_{n}$, with $D_{3}$ provided in the text and the theoretical gaussian expression in solid line by Gorbunova et al. (2021). (b) Similar representation of the longitudinal correlation function $C_{\text{\tiny{long}}}^{(4),\nu}(\tau,k_{n})$ (91) as in (a) but for the Gaussian model $\widehat{u}_{\text{\tiny{long}}}^{(4),\nu}(t,k_{n})$. Solid lines represent theoretical predictions while dot-lines are numerical estimations for the same wavenumbers $k_{n}$ as (a). The characteristic large time scale $T_{E}=1.43$ is defined by $T_{E}:=\sqrt{3}\frac{L_{\text{\tiny{int}}}}{\sigma^{\nu}}$ with $L_{\text{\tiny{int}}}$ being the integral time scale defined in (16) Inset: Re-scaled correlation functions in semi-log scale for the same wave numbers $k_{n}$, with the same $D_{3}$ as for DNS and provided in the text. (c) Numerical estimation of the characteristic time scale $T_{k_{n}}$ of the time correlation functions displayed in (a) and (b). We superimpose the numerical estimation of $T_{k_{n}}$ with a dashed black line given by the large $k$ asymptotic $(D_{3}k)^{-1}$ of the theoretical expression (53), with relevant free parameters provided in the text. Inset: re-scaled estimated $T_{k_{n}}$ by $(D_{3}k)^{-1}$, the asymptotical behavior of $T_{k}$ at large $k$.
@@
We show in Fig. 2(a) the results of our numerical estimation of the time correlation function of the DNS, $C_{\text{\tiny{long}}}^{\nu}(\tau,k_{n})$ (91) for several wave-number amplitudes $k_{n}$ and as a function of the time-lag $\tau$. At vanishing time lag $\tau=0$, all curves coincide with unity by construction, and then decrease toward 0 (they decorrelate) as $\tau$ increases. The characteristic time scale $T_{k_{\mathbf{n}}}^{\text{\tiny{DNS}}}$ of decorrelation is thus clearly a decreasing function of the wave number $k_{n}$. As has been observed by Gorbunova et al. (2021), the very shape of the correlation function is close to a Gaussian function, and so, to estimate the characteristic time scale $T_{k_{n}}^{\text{\tiny{DNS}}}$, we perform a similar Gaussian fit as by Gorbunova et al. (2021). In a straightforward manner, for each accessible $1/L_{\text{\tiny{tot}}}<k_{n}<512/L_{\text{\tiny{tot}}}$, we find the best possible $T_{k_{n}}^{\text{\tiny{DNS}}}$ such that the temporal correlation function $C_{\text{\tiny{long}}}^{\nu}(\tau,k_{n})$ (91) resembles the explicit Gaussian function $\exp(-(\tau/T_{k_{n}}^{\text{\tiny{DNS}}})^{2})$ and reproduce the result of this fitting procedure in Fig. 2(c) using an orange line. We can indeed observe that $T_{k_{n}}^{\text{\tiny{DNS}}}$ is a decreasing function of the wave number $k_{n}$, and follows a power-law, as has clearly been evidenced by the linear trends in this doubly logarithmic representation. We can also identify two ranges, below and above $k_{n}L_{\text{\tiny{tot}}}\approx 10$. This transition has already been observed in several studies (Kaneda et al., 1999; Favier et al., 2010; Gorbunova et al., 2021) and remains unexplained, as far as we know. We believe that at large scales, i.e. at low wave numbers, the flow is sensitive to boundary conditions and thus cannot be considered universal. At higher wave numbers $k_{n}L_{\text{\tiny{tot}}}$ greater than 10, the power decrease is consistent with the expected behavior given in (53) with $\beta=1/2$ (governing the power-law exponent) and $D_{3}=3.62$. We superimpose with a dashed black line the proposed function dependence of $T_{k_{n}}$ (53) in the limit of large $k_{n}$, with these aforementioned free parameters and clearly reproduce the decreasing behavior. To check whether such a parameterization of the characteristic time scale reproduces the behaviors of Fourier modes time correlation functions, we add in the inset of Fig. 2(a) a representation of $C_{\text{\tiny{long}}}^{\nu}(\tau,k_{n})$ (91) as a function of the re-scaled time variable $T_{k_{n}}^{\text{\tiny{DNS}}}$ in semi-log scale, and observe that all correlation functions collapse in a satisfactory manner on the expected theoretical curve.
@@
For comparison against our modeling approach, we consider the Gaussian velocity field $u^{{(4)},\nu}_{i}(t,\mathbf{x})$, i.e. for $N=4$. It is fully characterized in the statistically stationary regime by the covariance structure of its Fourier modes $\widehat{u}^{(4),\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})$ which is provided in (82). We have considered an evolution based on four layers $N=4$ that ensures a smooth behavior in time for a finite viscosity, as it is shown in (B). This choice is dictated by the necessity of ensuring smooth behavior in time of the velocity field, corresponding to bounded velocity time-derivatives (in this case, $\partial_{t}u^{(N),\nu}_{i}(t,\mathbf{x})$ have a finite variance for all integer $N>1$), as is observed in DNS. As we explain in Chapter 4, each layer needs to be stored in memory, which can become prohibitive given the size of the boxes, which corresponds to three-components of size $1024^{3}$ collocation points for each layer. Besides considering $H=1/3$, the free parameters $D_{2}$, $L$ and $\eta_{d}$ entering in the functional form of the longitudinal PSD (51) have the same values as the ones chosen to describe the spatial structure of the velocity field, as they are given in Section 2.2. The remaining free parameters $D_{3}$ and $\beta$ that appear in the characteristic correlation time scale $T_{k}$ of Fourier modes (53) are the ones that we have estimated before, corresponding to $\beta=1/2$ and $D_{3}=3.62$. We would like to recall that the value of $D_{3}$ is not universal because it is related, when $\beta=1/2$, to the standard deviation of the velocity. An efficient and exact in distribution numerical scheme is proposed in Section 4 in order to give instances in space and time of the model field $u^{(N),\nu}_{i}(t,\mathbf{x})$ for all integer $N>1$. Overall, the simulation of the vector field over more than five thousands time steps took approximatively two days using 32 cpus.
@@
As we did for the DNS velocity field, we display in Fig. 2(b) the temporal longitudinal correlation functions $C_{\text{\tiny{long}}}^{(N),\nu}(\tau,k_{n})$ (91) for the modeled Gaussian velocity field $\mathbf{u}^{(4),\nu}(t,\mathbf{x})$. Although being known to be different from the DNS in a theoretical point of view and provided in (82) (once recasted into the longitudinal context (93)), the estimations of $C_{\text{\tiny{long}}}^{(4),\nu}(\tau,k_{n})$ for various wave numbers $k_{n}$ displays similar trends as for the DNS. In this case, the collapse is not perfect because we are considering only the longitudinal correlation functions, which necessitate a transformation given in (93)) when going from the three dimensional formulation to the one dimensional setting. We display in Fig. 2(c) the time $T_{k}$ as a function of $k$, for the model for which Fourier mode temporal correlation is prescribed using blue symbols, and the measured Fourier mode correlation time for DNS in orange. We observe a very satisfactory collapse of the estimations of $T_{k}$ from DNS and the model. We also superimpose in dashed line the asymptotical limit $\left(D_{3}k_{n}\right)^{-1}$ using the form provided in (53). In the inset, we display the re-scaled correlation functions by the asymptotic behavior of the imposed characteristic time scale $\left(D_{3}k_{n}\right)^{-1}$ (53) at large $k_{n}$.
@@
Figure 3: Direct comparison of the temporal structure of the DNS and modeled velocity fields, and estimation of PSDs and second-order structure functions. (a) Spatio-temporal representation of the DNS velocity field $|\mathbf{u}^{\text{\tiny{DNS}},\nu}(t,\mathbf{x})|$ along the spatial line $\mathbf{x}\in([-\pi,\pi],0,0)$ and across time $t\in[0,5027\Delta t]$ where the time stepping $\Delta t$ is provided by the Hopkins database. We use the same characteristic large time scale $T_{E}^{\text{\tiny{DNS}}}$ as in Fig. 2 to adimensionalize time. (b) Similar representation as in panel (a) but for the model $|\mathbf{u}^{(4),\nu}(t,\mathbf{x})|$ with a characteristic large time scale $T_{E}$ of the order of $T_{E}^{\text{\tiny{DNS}}}$. We use the same colorbar for panels (a) and (b). (c) Estimation of the temporal power spectral density $E_{\nu}^{\text{\tiny{T}}}(\omega)$ (43) obtained as the variance of the temporal Fourier modes (see text). We use orange for DNS and blue for the model. The dashed black line corresponds to the power-law given in (B) without any fitting parameters. (d) Estimation of the second-order temporal structure function $S_{2}^{\text{\tiny{T}},\nu}(\tau)$ (given by (2.4) in the inviscid limit), same colors as in (c). Dashed black lines corresponds to two times the variance at large time lags, to (B) in the inertial range, and to (B) in the dissipative range.
@@
Our statistical estimation of the temporal power spectrum density $E_{\nu}^{\text{\tiny{T}}}(\omega)$ defined in (43) is shown in Fig. 3(c). To do so, we performed a discrete one-dimensional Fourier transform in time of the profile of the scalar product of velocity fields at a fixed position in space, as required in the definition (43). The obtained evolution in time is not periodic, and in doing so we introduce a spurious discontinuity in the picture. We have checked that windowing this profile aimed at minimizing the consequences of this discontinuity is not altering the observed behaviors in the inertial range of scales, but is slightly changing the results in the respective dissipative range (data not shown). As a result, the temporal spectra correspond to the variance of the (temporal) Fourier modes at each accessible frequency. The variances of the (temporal) Fourier modes of the velocity fields $\mathbf{u}^{\text{\tiny{DNS}},\nu}(t,\mathbf{x})$ and $\mathbf{u}^{(N),\nu}(t,\mathbf{x})$, obtained of the same statistical sample previously described, as functions of the frequency, are displayed doubly logarithmically in Fig. 3(c). We can see that with previously given free parameters, the temporal spectra superimpose in a very satisfactory manner in the inertial range, and follow a typical $\omega^{-5/3}$ power-law decrease.
@@
We finish with the estimation of the second-order temporal structure function $S_{2}^{\text{\tiny{T}},\nu}(\tau):=\mathbb{E}\left[\left(\mathbf{u}^{\nu}\left(t+\tau,\mathbf{x}\right)-\mathbf{u}^{\nu}\left(t,\mathbf{x}\right)\right)^{2}\right]$, as defined in (2.4) in the vanishing viscosity limit, and display our results in Fig. 3(d). The trends are consistent with the observations made for the temporal spectra. Both DNS and model velocity fields exhibit a power-law behavior both in the inertial and the dissipative ranges, with respective power-law exponents indicated on the figure. To see this, we superimpose the prediction made in (B) with a black dashed line. Note that in the inertial range of scales, predictions depend only weakly on the value of the number of layers N to define the model.
@@
To conclude this comparison, let us mention that additional comparisons could be done. In particular, we could also display the full spatio-temporal correlation function, as it has been done by Clark Di Leoni et al. (2015), and the frequency-wave spectrum as estimated by de Kat & Ganapathisubramani (2015), for both the DNS and the model. We are keeping this for further investigations. We would like also to mention that while the statistical structure of the DNS is fairly satisfactory (at least up to the aforementioned second-order statistical characterization), our model does not capture in a fully appropriate way the advection of the eddies by some large scale flow, as it is clear from the inspection of Fig. 3(a) and Fig. 3(b). To our knowledge, there is no model capable of doing this except some recent propositions by Armstrong & Vicol (2025), where the temporal structure of the advecting velocity field, although simplistic in terms of regularity, is defined recursively through scales using an inverse Lagrangian flow map at each scale of the construction. We will return to these possible perspectives in the final Section.